# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

### v1

@@ II.2.4 LMPW25 second-order classical algorithm
II.2.5 LMPW25 second-order hybrid quantum-classical algorithm
Finally, let us recall a second-order HQC algorithm for constrained energy minimization, as proposed in [16, Section VI]. A key observation, originally made in the context of a general-purpose natural gradient algorithm for quantum Boltzmann machines [81, Theorem 2], is that the Hessian matrix in (24) admits the following alternative expression [16, Lemma 6]:
@@ II.2.5 LMPW25 second-order hybrid quantum-classical algorithm
and $p(t)$ is the following high-peak tent probability density defined for $t\in\mathbb{R}$ [85, Eq. (9)]:
@@
Let us now write down the explicit steps of the LMPW25 second-order HQC algorithm for constrained energy minimization:
@@
where $C^{(j)}\equiv(C_{1}^{(j)},\ldots,C_{c}^{(j)})$. This simplification has the effect of reducing a multi-site Hamiltonian evolution, acting on all systems, to single-site evolutions and can significantly reduce the circuit depth. See Appendix B for a proof of (39).
@@ II.3 Quantum Heisenberg models
As mentioned in Section I.2, quantum Heisenberg models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38]. They are key examples considered for our simulations of the LMPW25 algorithms, the results of which we report in Section VI. A general form for a quantum Heisenberg model involves qubits placed at the vertices of a graph $G$ with vertex set $V$ and edge set $E$ [37]. The Hamiltonian is defined in terms of the graph $G$ and a list $\vec{J}\coloneqq(J_{i,j})_{\left\{i,j\right\}\in E}$ of coupling coefficients as follows:
@@
Special cases of the general model presented in (40)–(46) include the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions. In the case of a one-dimensional model, the qubits are arranged on a line with edges placed between nearest neighbors, for the nearest neighbor case, and edges additionally placed between next-to-nearest neighbors, for the next-to-nearest neighbor case. To simplify the model further, we set the coupling coefficients to be a fixed value $J$ between nearest neighbors, and for next-to-nearest neighbors, we set the coupling coefficients to be fixed at a constant value, say, $\lambda J$, where $\lambda\in[0,1]$. For a two-dimensional model, the main difference is that the qubits are arranged on a square lattice, and we take similar conventions, as mentioned above, for the nearest-neighbor and next-to-nearest neighbor cases.
@@ III Design of Hamiltonians, ground states, and thermal states
In this section, we offer alternative interpretations and applications of the constrained energy minimization and constrained free energy minimization problems in (12) and (13), respectively. Namely, we show how applying the LMPW25 algorithms to these problems leads to the design of Hamiltonians with ground and thermal states satisfying fixed constraints on the expectation values of certain observables. In this way, the LMPW25 algorithms could be useful for the design of molecules and materials satisfying desired properties. Let us note that the design of ground states has been considered previously in the context of the density matrix renormalization group, in the special case that there is a single chemical potential that constrains a single conserved charge of the system [89, 90, 91].
@@ III.1 Design of ground states
The first equality in (71) follows by introducing $\mu$ as a vector of Lagrange multipliers to enforce the constraints in (69). The equality in (72) follows by moving the supremum to the outside, and the equality in (73) follows by simple algebra. The penultimate equality follows from the Sion minimax theorem for extended reals [92, Theorem 2.11]. The final equality follows by bringing in the minimum over all density matrices and recognizing that
@@
In this way, by solving the constrained energy minimization problem in (70) and its dual in (75), we determine an optimal choice $\mu^{*}$ of the controllable parameters for the Hamiltonian $H-\mu^{*}\cdot Q$, such that there exists a state $\rho^{*}$ in the ground space of $H-\mu^{*}\cdot Q$ satisfying the constraints in (70). As such, by solving the primal and dual optimization problems, we can design Hamiltonians with a ground state satisfying desirable properties, i.e., the constraints in (69). The LMPW25 algorithms solve the constrained energy minimization problem approximately, by instead solving a free-energy minimization problem at low temperature, as reviewed in Section II.2, and preparing a low-temperature thermal state close to the ground state. After the LMPW25 HQC algorithm prepares this state, other observables could then be measured on it, in order to learn various properties of it.
@@ III.2 Design of thermal states
In this way, by solving the constrained free energy minimization problem in (78) and its dual in (91), we determine the optimal choice $\mu^{*}$ of the controllable parameters for the Hamiltonian $H-\mu^{*}\cdot Q$, such that $\rho_{T}(\mu^{*})$ is a temperature-$T$ thermal state of $H-\mu^{*}\cdot Q$ satisfying the constraints in (70). As such, by solving this dual optimization problem, we can design Hamiltonians with thermal states satisfying desirable properties, i.e., the constraints in (69). The LMPW25 algorithms solve this problem to arbitrary accuracy and thus can be used for the design of thermal states. If one is exclusively interested in the design of thermal states, note that it is not necessary for the algorithms to output the final value of the free energy, but they instead can exclusively perform the gradient update steps in Algorithms 14 until the norm of the gradient is sufficiently small.
@@ IV Stabilizer thermodynamic systems
where $i_{\ell}\in\{0,1,2,3\}$ and we have employed the notation introduced in (64). In this latter case, if all of the expectations $\left\langle Q_{i_{1},\ldots,i_{k}}\right\rangle_{\overline{\rho}}=r_{i_{1},\ldots,i_{k}}$ are fixed, where $\overline{\rho}$ is a state in the ground space of $H$ and $r_{i_{1},\ldots,i_{k}}\in\mathbb{R}$ for all $i_{1},\ldots,i_{k}\in\{0,1,2,3\}$, then the state of the $k$ logical qubits is completely specified after doing so (see (64)). This case leads to a particular way of using the LMPW25 HQC algorithms for encoding quantum information into the ground space of $H$, which we discuss in more detail in Section V.
@@ V Encoding quantum information into stabilizer thermodynamic systems
The LMPW25 first- and second-order HQC algorithms reviewed in Sections II.2.3 and II.2.5 can alternatively serve as methods for encoding quantum information into stabilizer thermodynamic systems. To see this, consider a stabilizer thermodynamic system corresponding to a stabilizer code encoding $k$ logical qubits into $n$ physical qubits, with a stabilizer $\mathcal{S}$ and a set $\mathcal{L}$ of logical operators. Suppose that we would like to encode a particular state of the $k$ logical qubits. Recall from (64) that such a state is specified by $4^{k}$ Pauli coefficients in the tuple $\left(r_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$. Then, after defining the Hamiltonian $H$ as in (97), with $\gamma_{i}=1$ for all $i\in[n-k]$, we set up a constrained energy minimization of the form in (12), as follows:
@@
By transforming this constrained energy minimization problem to a constrained free-energy minimization problem at temperature $T\geq 0$, as reviewed in Section II.2.1, the LMPW25 algorithms converge to an approximate solution of (108) for a sufficiently low temperature $T$. Furthermore, the HQC algorithms at the final step generate a parameterized thermal state of the following form:
@@
where the chemical potentials in $\left(\mu_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$ ensure that the constraints in (108) are satisfied. Thus, in this way, the LMPW25 HQC algorithms can be viewed as methods for approximately encoding logical qubit states into stabilizer thermodynamic systems, whenever low-temperature thermal state preparation methods are experimentally accessible.
@@ V.1 Warm-starting the encoding of a single qubit
The LMPW25 HQC algorithms start from the initial value $\mu=\left(0,\ldots,0\right)$, as seen in Algorithms 2 and 4. Such an initial starting point is used whenever there is no information about the solution to the problem at hand. However, in the case of encoding a single logical qubit into multiple physical qubits, we can provide a much better starting point for the algorithms, which speeds up the convergence significantly.
@@
The coordinates specified by $r$ are known as mixture coordinates, and those specified by $\mu$ and $\beta$ are known as exponential coordinates [94, Section 8.2]. The following theorem, proven in Appendix E, establishes a mapping from mixture coordinates to exponential coordinates for a single-qubit state:
@@
Equipped with Theorem 7, we can initialize the LMPW25 HQC algorithms with the following state:
@@
where $H$ is the stabilizer Hamiltonian for the code (as in (97), with $\gamma_{i}=1$ for all $i\in[n-k]$), $\overline{X}$, $\overline{Y}$, and $\overline{Z}$ are the logical operators for the encoded logical qubit, and $\mu$ and $\beta$ are chosen as in (113)–(114). The extra factor of $T$ multiplying $\beta$ in (116) ensures that the constraints in (108) are satisfied identically. In our simulations, we find that this initial starting point leads to immediate convergence of the LMPW25 HQC algorithms. In fact, the state in (116) is actually optimal, as a special case of the following more general theorem, which is proven in Appendix F:
@@ VI Simulation results
Figure 1: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the task of constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms require more iterations to converge due to sampling noise inherent in them.
In this section, we report the results of simulating the LMPW25 algorithms on several models of interest in quantum thermodynamics and the design of thermal states. Here, we focus on the one- and two-dimensional quantum Heisenberg models with a Hamiltonian having nearest- and next-to-nearest-neighbor interactions and the conserved, non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions (see Section II.3). We also report the results of simulations of a stabilizer thermodynamic system built from the perfect one-to-five qubit quantum error-correcting code.
@@
$\left|E-\tilde{E}_{m}\right|+\left\|\tilde{\nabla}_{m}\right\|,$ (122)
@@
in order to track the performance of the algorithms. In (122), $E$ is the true minimum energy output from a standard SDP solver and $\tilde{E}_{m}$ and $\tilde{\nabla}_{m}$ are the estimates of the energy and the gradient, respectively, at the $m$th step of the algorithm. If $\left|E-\tilde{E}_{m}\right|$ is small, then we are guaranteed that the algorithm has output a good approximation of the true energy at the $m$th step, and if $\left\|\tilde{\nabla}_{m}\right\|$ is small, then, according to (18), we are guaranteed that the algorithm has found a solution that satisfies the constraints in (12) approximately. In all of our figures, we plot the logarithm of the error metric in (122), in order to make it easier to visualize the performance of the algorithms.
Our first set of simulations involves the two-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem from Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be $0.5$, $0.5$, and $0.5$, respectively. Here we take the number of qubits to be fixed at six, arranged on a $2\times 3$ square lattice. Figure 1 compares the performance of all the LMPW25 algorithms, including the first- and second-order, classical and HQC algorithms, plotting the error metric in (122) versus the number of iterations. The figure demonstrates that all algorithms converge. The classical first-order algorithm employs Nesterov accelerated gradient, thus explaining the oscillatory behavior that occurs while converging. The classical second-order algorithm employs backtracking, as discussed after Algorithm 3. The HQC algorithms employ sampling, resulting in noisy estimates of the gradient and, for the second-order algorithm, the Hessian. We have plotted the average of five runs for both the first- and second-order HQC algorithms, along with the standard deviation shaded. We used roughly $10^{4}$ samples per iteration for the first-order HQC algorithm, and $10^{7}$ iterations for the second-order HQC algorithm, thus explaining why the first-order HQC algorithm is noisier than the second-order one. The figure demonstrates that incorporating second-derivative information, as done in the second-order algorithms, improves performance, i.e., requiring much fewer iterations to converge than do the first-order algorithms. However, as stated previously, the second-order algorithms require more computations per iteration than do the first-order algorithms, and so there is a trade-off present.
@@
Figure 2: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for all of the LMPW25 algorithms (1st- and 2nd-order, and classical and HQC) for the task of constrained energy minimization for a stabilizer thermodynamic system formed from the perfect five-qubit error-correcting code. The constraints on the logical operators $\overline{X}$, $\overline{Y}$, and $\overline{Z}$ were set to $0.2$, 0, and $0.5$, respectively. All of the algorithms converge, but the HQC algorithms require more iterations to converge due to sampling noise inherent in them. For this simulation, we did not warm-start the algorithm according to the recipe from Section V.1, but we instead started with $\mu_{x}=0$, $\mu_{y}=0$, and $\mu_{z}=0$.
@@
Figure 2 plots the performance of the LMPW25 algorithms for the above stabilizer thermodynamic system. Again the goal is to solve the constrained energy minimization problem from Section II.2.1. However, as mentioned in Section V, the LMPW25 HQC algorithms in this context can alternatively be understood as methods for encoding quantum information into an error-correcting code. Indeed, by setting the expectation values of the logical operators to be
@@
where $\left(r_{x},r_{y},r_{z}\right)\in\mathbb{R}^{3}$ is a vector in the Bloch ball, the LMPW25 HQC algorithms encode a logical qubit with this Bloch sphere representation. In our simulations, we choose
@@
Our final set of simulations involves the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem reviewed in Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. Figure 3 demonstrates the performance of the first-order LMPW25 classical algorithm as the number of qubits in the model increases (specifically, the figure plots results for three, five, and six qubits). The figure plots the error metric in (122) versus the number of iterations, demonstrating that every case converges to the true value as the number of iterations increases. The oscillatory behavior at the end of the algorithm is again due to using Nesterov accelerated gradient. The figure also demonstrates that the algorithm takes more iterations to converge to the true solution as the number of qubits increases.
Figure 3: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the task of constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW25 classical algorithm converges in all cases (3, 5, and 6 qubits), with larger systems taking more iterations to converge.
Figure 4 demonstrates the performance of the first-order HQC algorithm for the same setup. The algorithm converges in all cases, but it requires more iterations to do so than does the classical algorithm (about 10 times as many). As expected, this is due to the sampling noise that occurs when running the HQC algorithm. We set the number of shots used per iteration according to [16, Eq. (F17)], with $\varepsilon=10^{-4}$ and $\delta=10^{-2}$. In this case, the norm of the Pauli coefficients of the total magnetizations is $n$, where $n$ is the number of qubits.

### v2

@@ II.2.4 LMPW second-order classical algorithm
II.2.5 LMPW second-order hybrid quantum-classical algorithm
Finally, let us recall a second-order HQC algorithm for constrained energy minimization, as proposed in [16, Section VI]. A key observation, originally made in the context of a general-purpose natural gradient algorithm for quantum Boltzmann machines [82, Theorem 2], is that the Hessian matrix in (24) admits the following alternative expression [16, Lemma 6]:
@@ II.2.5 LMPW second-order hybrid quantum-classical algorithm
and $p(t)$ is the following high-peak tent probability density defined for $t\in\mathbb{R}$ [86, Eq. (9)]:
@@
Let us now write down the explicit steps of the LMPW second-order HQC algorithm for constrained energy minimization:
@@
where $C^{(j)}\equiv(C_{1}^{(j)},\ldots,C_{c}^{(j)})$. This simplification has the effect of reducing a multi-site Hamiltonian evolution, acting on all systems, to single-site evolutions, which act only on single systems, and can significantly reduce the circuit depth. See Appendix B for a proof of (39).
@@ II.3 Quantum Heisenberg models
As mentioned in Section I.2, quantum Heisenberg models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38]. They are key examples considered for our simulations of the LMPW algorithms, the results of which we report in Section VI. A general form for a quantum Heisenberg model involves qubits placed at the vertices of a graph $G$ with vertex set $V$ and edge set $E$ [37]. The Hamiltonian is defined in terms of the graph $G$ and a list $\vec{J}\coloneqq(J_{i,j})_{\left\{i,j\right\}\in E}$ of coupling coefficients as follows:
@@
Special cases of the general model presented in (40)–(46) include the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions. In the case of a one-dimensional model, the qubits are arranged on a line with edges placed between nearest neighbors, for the nearest neighbor case, and edges additionally placed between next-to-nearest neighbors, for the next-to-nearest neighbor case. To simplify the model further, we set the coupling coefficients to be a fixed value $J$ between nearest neighbors, and for next-to-nearest neighbors, we set the coupling coefficients to be fixed at a constant value, say, $\lambda J$, where $\lambda\in[0,1]$. For a two-dimensional model, the main difference is that the qubits are arranged on a square lattice, and we take similar conventions, as mentioned above, for the nearest-neighbor and next-to-nearest neighbor cases. See Figure 1 for a depiction of these models.
Figure 1: Depiction of one- and two-dimensional Heisenberg models. (Top) Qubits of the one-dimensional model are arranged on a line with edges placed between nearest neighbors (black lines), and edges additionally placed between next-to-nearest neighbors (blue curves). (Bottom) Qubits of the two-dimensional model are arranged on a square lattice, with nearest neighbor and next-to-nearest neighbor interactions shown in black and blue lines, respectively.
@@ III Design of Hamiltonians, ground states, and thermal states
In this section, we offer alternative interpretations and applications of the constrained energy minimization and constrained free energy minimization problems in (12) and (13), respectively. Namely, we show how applying the LMPW algorithms to these problems leads to the design of Hamiltonians with ground and thermal states satisfying fixed constraints on the expectation values of certain observables. In this way, the LMPW algorithms could be useful for the design of molecules and materials satisfying desired properties. Let us note that the design of ground states has been considered previously in the context of the density matrix renormalization group, in the special case that there is a single chemical potential that constrains a single conserved charge of the system [90, 91, 92].
@@ III.1 Design of ground states
The first equality in (71) follows by introducing $\mu$ as a vector of Lagrange multipliers to enforce the constraints in (69). The equality in (72) follows by moving the supremum to the outside, and the equality in (73) follows by simple algebra. The penultimate equality follows from the Sion minimax theorem for extended reals [93, Theorem 2.11]. The final equality follows by bringing in the minimum over all density matrices and recognizing that
@@
In this way, by solving the constrained energy minimization problem in (70) and its dual in (75), we determine an optimal choice $\mu^{*}$ of the controllable parameters for the Hamiltonian $H-\mu^{*}\cdot Q$, such that there exists a state $\rho^{*}$ in the ground space of $H-\mu^{*}\cdot Q$ satisfying the constraints in (70). As such, by solving the primal and dual optimization problems, we can design Hamiltonians with a ground state satisfying desirable properties, i.e., the constraints in (69). The LMPW algorithms solve the constrained energy minimization problem approximately, by instead solving a free-energy minimization problem at low temperature, as reviewed in Section II.2, and preparing a low-temperature thermal state close to the ground state. After the LMPW HQC algorithm prepares this state, other observables could then be measured on it, in order to learn various properties of it.
@@ III.2 Design of thermal states
In this way, by solving the constrained free energy minimization problem in (78) and its dual in (91), we determine the optimal choice $\mu^{*}$ of the controllable parameters for the Hamiltonian $H-\mu^{*}\cdot Q$, such that $\rho_{T}(\mu^{*})$ is a temperature-$T$ thermal state of $H-\mu^{*}\cdot Q$ satisfying the constraints in (70). As such, by solving this dual optimization problem, we can design Hamiltonians with thermal states satisfying desirable properties, i.e., the constraints in (69). The LMPW algorithms solve this problem to arbitrary accuracy and thus can be used for the design of thermal states. If one is exclusively interested in the design of thermal states, note that it is not necessary for the algorithms to output the final value of the free energy, but they instead can exclusively perform the gradient update steps in Algorithms 14 until the norm of the gradient is sufficiently small.
@@ IV Stabilizer thermodynamic systems
where $i_{\ell}\in\{0,1,2,3\}$ and we have employed the notation introduced in (64). In this latter case, if all of the expectations $\left\langle Q_{i_{1},\ldots,i_{k}}\right\rangle_{\overline{\rho}}=r_{i_{1},\ldots,i_{k}}$ are fixed, where $\overline{\rho}$ is a state in the ground space of $H$ and $r_{i_{1},\ldots,i_{k}}\in\mathbb{R}$ for all $i_{1},\ldots,i_{k}\in\{0,1,2,3\}$, then the state of the $k$ logical qubits is completely specified after doing so (see (64)). This case leads to a particular way of using the LMPW HQC algorithms for encoding quantum information into the ground space of $H$, which we discuss in more detail in Section V.
@@ V Encoding quantum information into stabilizer thermodynamic systems
The LMPW first- and second-order HQC algorithms reviewed in Sections II.2.3 and II.2.5 can alternatively serve as methods for encoding quantum information into stabilizer thermodynamic systems. To see this, consider a stabilizer thermodynamic system corresponding to a stabilizer code encoding $k$ logical qubits into $n$ physical qubits, with a stabilizer $\mathcal{S}$ and a set $\mathcal{L}$ of logical operators. Suppose that we would like to encode a particular state of the $k$ logical qubits. Recall from (64) that such a state is specified by $4^{k}$ Pauli coefficients in the tuple $\left(r_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$. Then, after defining the Hamiltonian $H$ as in (97), with $\gamma_{i}=1$ for all $i\in[n-k]$, we set up a constrained energy minimization of the form in (12), as follows:
@@
By transforming this constrained energy minimization problem to a constrained free-energy minimization problem at temperature $T\geq 0$, as reviewed in Section II.2.1, the LMPW algorithms converge to an approximate solution of (108) for a sufficiently low temperature $T$. Furthermore, the HQC algorithms at the final step generate a parameterized thermal state of the following form:
@@
where the chemical potentials in $\left(\mu_{i_{1},\ldots,i_{k}}\right)_{i_{1},\ldots,i_{k}\in\{0,1,2,3\}}$ ensure that the constraints in (108) are satisfied. Thus, in this way, the LMPW HQC algorithms can be viewed as methods for approximately encoding logical qubit states into stabilizer thermodynamic systems, whenever low-temperature thermal state preparation methods are experimentally accessible.
@@ V.1 Warm-starting the encoding of a single qubit
The LMPW HQC algorithms start from the initial value $\mu=\left(0,\ldots,0\right)$, as seen in Algorithms 2 and 4. Such an initial starting point is used whenever there is no information about the solution to the problem at hand. However, in the case of encoding a single logical qubit into multiple physical qubits, we can provide a much better starting point for the algorithms, which speeds up the convergence significantly.
@@
The coordinates specified by $r$ are known as mixture coordinates, and those specified by $\mu$ and $\beta$ are known as exponential coordinates [95, Section 8.2]. The following theorem, proven in Appendix E, establishes a mapping from mixture coordinates to exponential coordinates for a single-qubit state:
@@
Equipped with Theorem 7, we can initialize the LMPW HQC algorithms with the following state:
@@
where $H$ is the stabilizer Hamiltonian for the code (as in (97), with $\gamma_{i}=1$ for all $i\in[n-k]$), $\overline{X}$, $\overline{Y}$, and $\overline{Z}$ are the logical operators for the encoded logical qubit, and $\mu$ and $\beta$ are chosen as in (113)–(114). The extra factor of $T$ multiplying $\beta$ in (116) ensures that the constraints in (108) are satisfied identically. In our simulations, we find that this initial starting point leads to immediate convergence of the LMPW HQC algorithms. In fact, the state in (116) is actually optimal, as a special case of the following more general theorem, which is proven in Appendix F:
@@ VI Simulation results
Figure 2: The figure depicts the logarithm of the error metric in (122) versus the logarithm of the number of iterations, for the task of constrained energy minimization for the two-dimensional, four-qubit quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them.
In this section, we report the results of simulating the LMPW algorithms on several models of interest in quantum thermodynamics and the design of thermal states. Here, we focus on the one- and two-dimensional quantum Heisenberg models with a Hamiltonian having nearest- and next-to-nearest-neighbor interactions and the conserved, non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions (see Section II.3). We also report the results of simulations of a stabilizer thermodynamic system built from the perfect one-to-five qubit quantum error-correcting code.
@@
$\left|E^{\star}-\tilde{E}_{m}\right|+\left\|\tilde{\nabla}_{m}\right\|,$ (122)
@@
in order to track the performance of the algorithms. In (122), $E^{\star}$ is the true minimum energy output from a standard SDP solver and $\tilde{E}_{m}$ and $\tilde{\nabla}_{m}$ are the estimates of the energy and the gradient, respectively, at the $m$th step of the algorithm. If $\left|E^{\star}-\tilde{E}_{m}\right|$ is small, then we are guaranteed that the algorithm has output a good approximation of the true energy at the $m$th step, and if $\left\|\tilde{\nabla}_{m}\right\|$ is small, then, according to (18), we are guaranteed that the algorithm has found a solution that satisfies the constraints in (12) approximately. In all of our figures, we plot the logarithm of the error metric in (122), in order to make it easier to visualize the performance of the algorithms. In addition to using this error metric for plotting purposes, we also used the gradient norm $\left\|\tilde{\nabla}_{m}\right\|$ as the stopping criterion to determine algorithm convergence, terminating the optimization once it fell below a specified threshold.
@@
Figure 3: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for all of the LMPW algorithms (1st- and 2nd-order, and classical and HQC) for the task of constrained energy minimization for a stabilizer thermodynamic system formed from the perfect five-qubit error-correcting code. The constraints on the logical operators $\overline{X}$, $\overline{Y}$, and $\overline{Z}$ were set to $0.2$, 0, and $0.5$, respectively. All of the algorithms converge, but the HQC algorithms, shown as the average over five independent runs with shaded regions denoting one standard deviation, require more iterations to converge due to sampling noise inherent in them. For this simulation, we did not warm-start the algorithm according to the recipe from Section V.1, but we instead started with $\mu_{x}=1$, $\mu_{y}=1$, and $\mu_{z}=1$.
Our first set of simulations involves the two-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and with the charges set to the total magnetizations in the $x$, $y$, and $z$ directions. The goal is to solve the constrained energy minimization problem from Section II.2.1, with the expectation values on the total magnetizations in the $x$, $y$, and $z$ directions set to be $1$, 0, and $1$, respectively. Here we take the number of qubits to be fixed at four, arranged on a $2\times 2$ square lattice. Figure 2 compares the performance of all the LMPW algorithms, including the first- and second-order, classical and HQC algorithms, plotting the error metric in (122) versus the number of iterations. For the first- and second-order HQC algorithms, the average of five runs, along with one standard deviation shaded, is plotted. The figure demonstrates that all algorithms converge.
The classical and HQC second-order algorithms employ backtracking, as discussed after Algorithm 3. The HQC algorithms employ sampling, resulting in noisy estimates of the gradient and, for the second-order algorithm, the Hessian. We used approximately $10^{4}$ samples per iteration to estimate expectation values of observables in both the first-order and second-order HQC algorithms, and $10^{7}$ samples per iteration were employed to estimate the Hessian for the second-order HQC algorithm. To reduce the computational cost of Hessian estimation, we parallelized the simulation across multiple CPU cores using process-level parallelism. Each process evaluates a subset of Hadamard-test circuits with a dedicated, single-threaded simulator to avoid resource contention. The circuits are pre-transpiled with placeholders for the time-evolution unitaries, which are dynamically inserted at runtime using cached matrix exponentials. This strategy significantly accelerates the Hessian estimation step while keeping memory usage and inter-process contention low. As shown in the figure, the second-order HQC exhibits a higher variance across independent runs, primarily due to the statistical and numerical challenges involved in estimating and inverting the Hessian. In particular, the inverse Hessian estimate is significantly more sensitive to noise and often suffers from numerical instabilities, which can amplify fluctuations in the update direction. To mitigate the computational cost of Hessian estimation, we adopted a looser convergence threshold for the gradient norm in the second-order HQC algorithm. This explains the comparatively higher error metric reached by the second-order HQC algorithm: unlike the other algorithms, whose faster runtime permitted more iterations and finer convergence, the second-order HQC was terminated earlier to maintain computational feasibility.
The figure clearly illustrates that incorporating second-derivative information, as done in the second-order algorithms, improves performance, i.e., requiring much fewer iterations to reach convergence compared to the first-order methods. However, as stated previously, the second-order algorithms require more computations per iteration than the first-order algorithms, and thus a trade-off is present.
@@
Figure 4: The figure depicts the logarithm of the error metric in (122) versus the number of iterations, for the task of constrained energy minimization for the one-dimensional quantum Heisenberg model with nearest- and next-to-nearest-neighbor interactions and constraints on the total magnetizations in the $x$, $y$, and $z$ directions set to be 1, 0, and 1, respectively. The LMPW classical algorithm converges in all cases (3, 4, 5, and 6 qubits), with larger systems taking more iterations to converge.
Figure 3 plots the performance of the LMPW algorithms for the above stabilizer thermodynamic system. Again the goal is to solve the constrained energy minimization problem from Section II.2.1. However, as mentioned in Section V, the LMPW HQC algorithms in this context can alternatively be understood as methods for encoding quantum information into an error-correcting code. Indeed, by setting the expectation values of the logical operators to be
@@
where $\left(r_{x},r_{y},r_{z}\right)\in\mathbb{R}^{3}$ is a vector in the Bloch ball, the LMPW HQC algorithms encode a logical qubit with this Bloch sphere representation. In our simulations, we choose